# Dwarf Galaxies in the MATLAS Survey: Hubble Space Telescope Observations of Nuclear Star Clusters

### v1

@@ 9 Summary and conclusions
In this work, we investigated the formation and evolution of NSCs by studying their structural and photometric properties in a dwarf galaxy sample biased towards low-surface-brightness and diffuse –UDG-like– galaxies located beyond the LV and outside the environment of galaxy clusters. We made use of HST follow-up observations of 79 MATLAS galaxies with the ACS using the F606W and F814W filters. This sample is composed of 41 nucleated dwarfs, 13 of which are newly identified thanks to HST high spatial resolution images, and two have a double nucleus.
@@
We modeled the NSCs with a Sérsic profile and obtained a reliable result for 31 of them; the derived parameters are provided in Table 1. We compared their structural properties and photometry to NSCs in nearby field LTGs, and in dwarf galaxies from the LV, as well as from the Virgo and Fornax galaxy clusters. The Re,NSC of the MATLAS NSCs does not tend to be larger as they get brighter and more massive. Moreover, while we do not see an influence of the absolute magnitude on the ellipticity, the Sérsic index of the MATLAS NSCs gets larger with M, as observed in the Virgo cluster.
@@
Looking at the relation between the NSCs and their host galaxy, we find that a non negligible fraction of the MATLAS NSCs tend to be brighter and more massive than NSCs in typical dwarf galaxies. This could be either due to the galaxies being ongoing tidal disruption or to an enhanced growth of NSCs in UDGs, producing more massive NSCs.
@@
We derived the color profiles within 3Re,NSC of the modeled NSCs and remark that, while the low-mass NSCs tend to have an almost constant profile consistent with the migration scenario, the most massive NSCs have a steep slope, from an outer redder color to an inner bluer color, suggesting a change of stellar population, which might be caused by a contribution from in-situ star formation.
@@
Following the work of Marleau et al. (2024) on the GC systems of the HST MATLAS UDGs, we studied the distribution and properties of the GCs in the context of the migration scenario. We modeled the GCs in the MATLAS dwarfs up to 25 Mpc and find that , although with low statistics, the most massive GCs with M¿ 10M⊙ are distributed oppositely in the nucleated and non-nucleated dwarfs: the former seem to lack massive GCs in their central 0.5Re,gal region, while they are located within 0.5Re,gal for the latter. Comparing the Re – luminosity scaling relation of the GCs to the NSCs, we find a different relation for both types of star clusters, in agreement with Bekki et al. (2004).
@@
The most massive UCDs are believed to be the NSC remnants of tidally stripped nucleated dwarfs. In that context, we searched for UCD progenitors among the 41 NSCs with available photometry. We found that the NSCs appear to link the GC and UCD populations in the Re – luminosity scaling relation, suggesting that some of them could become UCDs if their dwarf host undergo tidal disruption. Looking for strongly nucleated dwarfs, we identified two UCD progenitor candidates, MATLAS-1539 and MATLAS-1577, that have a nuclear-to-total-luminosity fraction of 6.7% and 8.7%, respectively, and show signs of a possible tidal interaction with their assumed host ETG. Regarding the galaxies with a low nuclear-to-total-luminosity fraction, we found three tidally elongated UDGs, likely in the early stage of tidal disruption, that we added to our list of UCD progenitor candidates.
@@
Overall, the observed trends need to be confirmed by studying even larger samples of nucleated dwarfs using deep and high resolution imaging facility (e.g., the Euclid Wide Survey; Euclid Collaboration: Scaramella et al. 2022), combined to spectroscopic follow-up observations to trace the NSC star formation history.
@@ Acknowledgements.
The authors would like to thank the referee for the insightful comments, which helped to improve the manuscript. This research is based on observations from the NASA/ESA Hubble Space Telescope obtained at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. Support for Program number GO-16257 and GO-16711 was provided through a grant from the STScI under NASA contract NAS5-26555. This research was supported by the International Space Science Institute (ISSI) in Bern, through ISSI International Team project No. 534. This research made use of AutoProf, a package for galaxy image photometry (Stone et al. 2021). M.P. is supported by the Academy of Finland grant No. 347089. O.M. and N.H. are grateful to the Swiss National Science Foundation for financial support under the grant number PZ00P2_202104. P.R.D. gratefully acknowledges support from grant HST-GO-16257.002-A. S.P. acknowledges support from the Mid-career Researcher Program (No. RS-2023-00208957). S.L. acknowledges the support from the Sejong Science Fellowship Program by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. NRF-2021R1C1C2006790). RH acknowledges funding from the Italian INAF Large Grant 12 - 2022.References
- Albareti et al. (2017) Albareti, F. D., Allende Prieto, C., Almeida, A., et al. 2017, ApJS, 233, 25
- Alfaro-Cuello et al. (2019) Alfaro-Cuello, M., Kacharov, N., Neumayer, N., et al. 2019, ApJ, 886, 57
- Ann et al. (2015) Ann, H. B., Seo, M., & Ha, D. K. 2015, ApJS, 217, 27
- Antonini et al. (2015) Antonini, F., Barausse, E., & Silk, J. 2015, ApJ, 812, 72
- Antonini et al. (2012) Antonini, F., Capuzzo-Dolcetta, R., Mastrobuono-Battisti, A., & Merritt, D. 2012, ApJ, 750, 111
- Arca-Sedda & Capuzzo-Dolcetta (2014) Arca-Sedda, M. & Capuzzo-Dolcetta, R. 2014, ApJ, 785, 51
- Arca-Sedda et al. (2015) Arca-Sedda, M., Capuzzo-Dolcetta, R., Antonini, F., & Seth, A. 2015, ApJ, 806, 220
- Balcells et al. (2003) Balcells, M., Graham, A. W., Domínguez-Palmero, L., & Peletier, R. F. 2003, ApJ, 582, L79
- Balcells et al. (2007) Balcells, M., Graham, A. W., & Peletier, R. F. 2007, ApJ, 665, 1084
- Bassino et al. (1994) Bassino, L. P., Muzzio, J. C., & Rabolli, M. 1994, ApJ, 431, 634
- Bekki (2007) Bekki, K. 2007, PASA, 24, 77
- Bekki et al. (2001) Bekki, K., Couch, W. J., & Drinkwater, M. J. 2001, ApJ, 552, L105
- Bekki et al. (2004) Bekki, K., Couch, W. J., Drinkwater, M. J., & Shioya, Y. 2004, ApJ, 610, L13
- Bell et al. (2003) Bell, E. F., McIntosh, D. H., Katz, N., & Weinberg, M. D. 2003, ApJS, 149, 289
- Bertin & Arnouts (1996) Bertin, E. & Arnouts, S. 1996, A&AS, 117, 393
- Bílek et al. (2020) Bílek, M., Duc, P.-A., Cuillandre, J.-C., et al. 2020, MNRAS, 498, 2138
- Binggeli et al. (1985) Binggeli, B., Sandage, A., & Tammann, G. A. 1985, AJ, 90, 1681
- Blanton et al. (2017) Blanton, M. R., Bershady, M. A., Abolfathi, B., et al. 2017, AJ, 154, 28
- Bohlin (2016) Bohlin, R. C. 2016, AJ, 152, 60
- Böker et al. (2002) Böker, T., Laine, S., van der Marel, R. P., et al. 2002, AJ, 123, 1389
- Böker et al. (2004) Böker, T., Sarzi, M., McLaughlin, D. E., et al. 2004, AJ, 127, 105
- Bradley et al. (2020) Bradley, L., Sipőcz, B., Robitaille, T., et al. 2020, astropy/photutils: 1.0.0
- Buzzo et al. (2025) Buzzo, M. L., Forbes, D. A., Jarrett, T. H., et al. 2025, MNRAS, 536, 2536
- Cappellari et al. (2011) Cappellari, M., Emsellem, E., Krajnović, D., et al. 2011, MNRAS, 413, 813
- Carlsten et al. (2022) Carlsten, S. G., Greene, J. E., Beaton, R. L., & Greco, J. P. 2022, ApJ, 927, 44
- Carollo et al. (1997) Carollo, C. M., Stiavelli, M., de Zeeuw, P. T., & Mack, J. 1997, AJ, 114, 2366
- Carson et al. (2015) Carson, D. J., Barth, A. J., Seth, A. C., et al. 2015, AJ, 149, 170
- Chilingarian et al. (2011) Chilingarian, I. V., Mieske, S., Hilker, M., & Infante, L. 2011, MNRAS, 412, 1627
- Côté et al. (2006) Côté, P., Piatek, S., Ferrarese, L., et al. 2006, ApJS, 165, 57
- Danieli et al. (2022) Danieli, S., van Dokkum, P., Trujillo-Gomez, S., et al. 2022, ApJ, 927, L28
- De Bórtoli et al. (2020) De Bórtoli, B. J., Bassino, L. P., Caso, J. P., & Ennis, A. I. 2020, MNRAS, 492, 4313
- den Brok et al. (2014) den Brok, M., Peletier, R. F., Seth, A., et al. 2014, MNRAS, 445, 2385
- Drinkwater et al. (2000) Drinkwater, M. J., Jones, J. B., Gregg, M. D., & Phillipps, S. 2000, PASA, 17, 227
- Duc (2020) Duc, P.-A. 2020, arXiv e-prints, arXiv:2007.13874
- Duc et al. (2015) Duc, P.-A., Cuillandre, J.-C., Karabal, E., et al. 2015, MNRAS, 446, 120
- Erwin (2015) Erwin, P. 2015, ApJ, 799, 226
- Erwin & Gadotti (2012) Erwin, P. & Gadotti, D. A. 2012, Advances in Astronomy, 2012, 946368
- Euclid Collaboration: Scaramella et al. (2022) Euclid Collaboration: Scaramella, R., Amiaux, J., Mellier, Y., et al. 2022, A&A, 662, A112
- Evstigneeva et al. (2008) Evstigneeva, E. A., Drinkwater, M. J., Peng, C. Y., et al. 2008, AJ, 136, 461
- Evstigneeva et al. (2007) Evstigneeva, E. A., Gregg, M. D., Drinkwater, M. J., & Hilker, M. 2007, AJ, 133, 1722
- Faber et al. (1997) Faber, S. M., Tremaine, S., Ajhar, E. A., et al. 1997, AJ, 114, 1771
- Fahrion et al. (2024) Fahrion, K., Böker, T., Perna, M., et al. 2024, A&A, 687, A83
- Fahrion et al. (2022a) Fahrion, K., Bulichi, T.-E., Hilker, M., et al. 2022a, A&A, 667, A101
- Fahrion et al. (2022b) Fahrion, K., Leaman, R., Lyubenova, M., & van de Ven, G. 2022b, A&A, 658, A172
- Fahrion et al. (2021) Fahrion, K., Lyubenova, M., van de Ven, G., et al. 2021, A&A, 650, A137
- Faifer et al. (2017) Faifer, F. R., Escudero, C. G., Scalia, M. C., et al. 2017, A&A, 599, L8
- Ferguson & Binggeli (1994) Ferguson, H. C. & Binggeli, B. 1994, A&A Rev., 6, 67
- Ferrarese et al. (2012) Ferrarese, L., Côté, P., Cuillandre, J.-C., et al. 2012, ApJS, 200, 4
- Ferrarese et al. (2006) Ferrarese, L., Côté, P., Jordán, A., et al. 2006, ApJS, 164, 334
- Forbes & Gannon (2024) Forbes, D. A. & Gannon, J. 2024, MNRAS, 528, 608
- Freeman (1993) Freeman, K. C. 1993, in Astronomical Society of the Pacific Conference Series, Vol. 48, The Globular Cluster-Galaxy Connection, ed. G. H. Smith & J. P. Brodie, 608
- Gaia Collaboration et al. (2018) Gaia Collaboration, Brown, A. G. A., Vallenari, A., et al. 2018, A&A, 616, A1
- Georgiev & Böker (2014) Georgiev, I. Y. & Böker, T. 2014, MNRAS, 441, 3570
- Georgiev et al. (2016) Georgiev, I. Y., Böker, T., Leigh, N., Lützgendorf, N., & Neumayer, N. 2016, MNRAS, 457, 2122
- Goerdt et al. (2008) Goerdt, T., Moore, B., Kazantzidis, S., et al. 2008, MNRAS, 385, 2136
- Graham (2020) Graham, A. W. 2020, MNRAS, 492, 3263
- Graham & Guzmán (2003) Graham, A. W. & Guzmán, R. 2003, AJ, 125, 2936
- Graham & Spitler (2009) Graham, A. W. & Spitler, L. R. 2009, MNRAS, 397, 2148
- Guillard et al. (2016) Guillard, N., Emsellem, E., & Renaud, F. 2016, MNRAS, 461, 3620
- Haacke et al. (2025) Haacke, L., Forbes, D. A., Gannon, J. S., et al. 2025, MNRAS, 539, 674
- Habas et al. (2020) Habas, R., Marleau, F. R., Duc, P.-A., et al. 2020, MNRAS, 491, 1901
- Harris (1996) Harris, W. E. 1996, AJ, 112, 1487
- Heesters et al. (2023) Heesters, N., Müller, O., Marleau, F. R., et al. 2023, A&A, 676, A33
- Hernandez & Gilmore (1998) Hernandez, X. & Gilmore, G. 1998, MNRAS, 297, 517
- Hilker et al. (1999) Hilker, M., Infante, L., Vieira, G., Kissler-Patig, M., & Richtler, T. 1999, A&AS, 134, 75
- Hoyer et al. (2023) Hoyer, N., Neumayer, N., Seth, A. C., Georgiev, I. Y., & Greene, J. E. 2023, MNRAS, 520, 4664
- Janz et al. (2016) Janz, J., Norris, M. A., Forbes, D. A., et al. 2016, MNRAS, 456, 617
- Johnston et al. (2020) Johnston, E. J., Puzia, T. H., D’Ago, G., et al. 2020, MNRAS, 495, 2247
- Kacharov et al. (2022) Kacharov, N., Alfaro-Cuello, M., Neumayer, N., et al. 2022, ApJ, 939, 118
- Krist et al. (2011) Krist, J. E., Hook, R. N., & Stoehr, F. 2011, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 8127, Optical Modeling and Performance Predictions V, ed. M. A. Kahan, 81270J
- Kroupa (1998) Kroupa, P. 1998, MNRAS, 300, 200
- Lim et al. (2020) Lim, S., Côté, P., Peng, E. W., et al. 2020, ApJ, 899, 69
- Liu et al. (2020) Liu, C., Côté, P., Peng, E. W., et al. 2020, ApJS, 250, 17
- Loose et al. (1982) Loose, H. H., Kruegel, E., & Tutukov, A. 1982, A&A, 105, 342
- Lotz et al. (2001) Lotz, J. M., Telford, R., Ferguson, H. C., et al. 2001, ApJ, 552, 572
- Lyubenova & Tsatsi (2019) Lyubenova, M. & Tsatsi, A. 2019, A&A, 629, A44
- Lyubenova et al. (2013) Lyubenova, M., van den Bosch, R. C. E., Côté, P., et al. 2013, MNRAS, 431, 3364
- Madrid et al. (2010) Madrid, J. P., Graham, A. W., Harris, W. E., et al. 2010, ApJ, 722, 1707
- Magnier & Cuillandre (2004) Magnier, E. A. & Cuillandre, J. C. 2004, PASP, 116, 449
- Marleau et al. (2025) Marleau, F. R., Cuillandre, J. C., Cantiello, M., et al. 2025, A&A, 697, A12
- Marleau et al. (2024) Marleau, F. R., Duc, P.-A., Poulain, M., et al. 2024, A&A, 690, A339
- Marleau et al. (2021) Marleau, F. R., Habas, R., Poulain, M., et al. 2021, A&A, 654, A105
- Matthews et al. (1999) Matthews, L. D., Gallagher, III, J. S., Krist, J. E., et al. 1999, AJ, 118, 208
- Mayes et al. (2024) Mayes, R. J., Drinkwater, M. J., Pfeffer, J., & Baumgardt, H. 2024, MNRAS, 527, 4643
- Mayes et al. (2021) Mayes, R. J., Drinkwater, M. J., Pfeffer, J., et al. 2021, MNRAS, 501, 1852
- Mieske et al. (2002) Mieske, S., Hilker, M., & Infante, L. 2002, A&A, 383, 823
- Mieske et al. (2008) Mieske, S., Hilker, M., Jordán, A., et al. 2008, A&A, 487, 921
- Mieske et al. (2012) Mieske, S., Hilker, M., & Misgeld, I. 2012, A&A, 537, A3
- Misgeld & Hilker (2011) Misgeld, I. & Hilker, M. 2011, MNRAS, 414, 3699
- Müller et al. (2021) Müller, O., Durrell, P. R., Marleau, F. R., et al. 2021, ApJ, 923, 9
- Müller et al. (2020) Müller, O., Marleau, F. R., Duc, P.-A., et al. 2020, A&A, 640, A106
- Neumayer et al. (2020) Neumayer, N., Seth, A., & Böker, T. 2020, A&A Rev., 28, 4
- Norris et al. (2015) Norris, M. A., Escudero, C. G., Faifer, F. R., et al. 2015, MNRAS, 451, 3615
- Norris & Kannappan (2011) Norris, M. A. & Kannappan, S. J. 2011, MNRAS, 414, 739
- Norris et al. (2014) Norris, M. A., Kannappan, S. J., Forbes, D. A., et al. 2014, MNRAS, 443, 1151
- Oh & Lin (2000) Oh, K. S. & Lin, D. N. C. 2000, ApJ, 543, 620
- Ordenes-Briceño et al. (2018) Ordenes-Briceño, Y., Puzia, T. H., Eigenthaler, P., et al. 2018, ApJ, 860, 4
- Paudel et al. (2023) Paudel, S., Duc, P.-A., Lim, S., et al. 2023, MNRAS, 526, L136
- Paudel et al. (2010) Paudel, S., Lisker, T., & Janz, J. 2010, ApJ, 724, L64
- Paudel et al. (2011) Paudel, S., Lisker, T., & Kuntschner, H. 2011, MNRAS, 413, 1764
- Paudel & Yoon (2020) Paudel, S. & Yoon, S.-J. 2020, ApJL, 898, L47
- Peacock et al. (2010) Peacock, M. B., Maccarone, T. J., Knigge, C., et al. 2010, MNRAS, 402, 803
- Pechetti et al. (2020) Pechetti, R., Seth, A., Neumayer, N., et al. 2020, ApJ, 900, 32
- Peng et al. (2002) Peng, C. Y., Ho, L. C., Impey, C. D., & Rix, H.-W. 2002, AJ, 124, 266
- Peng et al. (2010) Peng, C. Y., Ho, L. C., Impey, C. D., & Rix, H.-W. 2010, AJ, 139, 2097
- Pfeffer et al. (2014) Pfeffer, J., Griffen, B. F., Baumgardt, H., & Hilker, M. 2014, MNRAS, 444, 3670
- Pfeffer et al. (2024) Pfeffer, J., Janssens, S. R., Buzzo, M. L., et al. 2024, MNRAS, 529, 4914
- Phillipps et al. (2001) Phillipps, S., Drinkwater, M. J., Gregg, M. D., & Jones, J. B. 2001, ApJ, 560, 201
- Poulain et al. (2021) Poulain, M., Marleau, F. R., Habas, R., et al. 2021, MNRAS, 506, 5494
- Poulain et al. (2022) Poulain, M., Marleau, F. R., Habas, R., et al. 2022, A&A, 659, A14
- Poulain et al. (2025) Poulain, M., Smith, R., Duc, P.-A., et al. 2025, Nature, 640, 902
- Reaves (1983) Reaves, G. 1983, ApJS, 53, 375
- Roediger & Courteau (2015) Roediger, J. C. & Courteau, S. 2015, MNRAS, 452, 3209
- Rossa et al. (2006) Rossa, J., van der Marel, R. P., Böker, T., et al. 2006, The Astronomical Journal, 132, 1074
- Saifollahi et al. (2021) Saifollahi, T., Janz, J., Peletier, R. F., et al. 2021, MNRAS, 504, 3580
- Sánchez-Janssen et al. (2019) Sánchez-Janssen, R., Côté, P., Ferrarese, L., et al. 2019, ApJ, 878, 18
- Sandage et al. (1985) Sandage, A., Binggeli, B., & Tammann, G. A. 1985, AJ, 90, 1759
- Schiavi et al. (2021) Schiavi, R., Capuzzo-Dolcetta, R., Georgiev, I. Y., Arca-Sedda, M., & Mastrobuono-Battisti, A. 2021, MNRAS, 503, 594
- Schlafly & Finkbeiner (2011) Schlafly, E. F. & Finkbeiner, D. P. 2011, ApJ, 737, 103
- Schweizer et al. (2018) Schweizer, F., Seitzer, P., Whitmore, B. C., Kelson, D. D., & Villanueva, E. V. 2018, ApJ, 853, 54
- Scott & Graham (2013) Scott, N. & Graham, A. W. 2013, ApJ, 763, 76
- Sérsic (1963) Sérsic, J. L. 1963, BAAA, 6, 41
- Seth et al. (2010) Seth, A. C., Cappellari, M., Neumayer, N., et al. 2010, ApJ, 714, 713
- Sills et al. (2019) Sills, A., Dalessandro, E., Cadelano, M., Alfaro-Cuello, M., & Kruijssen, J. M. D. 2019, MNRAS, 490, L67
- Sirianni et al. (2005) Sirianni, M., Jee, M. J., Benítez, N., et al. 2005, PASP, 117, 1049
- Smithsonian Astrophysical Observatory (2000) Smithsonian Astrophysical Observatory. 2000, SAOImage DS9: A utility for displaying astronomical images in the X11 window environment, Astrophysics Source Code Library, record ascl:0003.002
- Spengler et al. (2017) Spengler, C., Côté, P., Roediger, J., et al. 2017, ApJ, 849, 55
- Stone et al. (2021) Stone, C. J., Arora, N., Courteau, S., & Cuillandre, J.-C. 2021, MNRAS, 508, 1870
- Su et al. (2022) Su, A. H., Salo, H., Janz, J., Venhola, A., & Peletier, R. F. 2022, A&A, 664, A167
- Tremaine et al. (1975) Tremaine, S. D., Ostriker, J. P., & Spitzer, L., J. 1975, ApJ, 196, 407
- Turner et al. (2012) Turner, M. L., Côté, P., Ferrarese, L., et al. 2012, ApJS, 203, 5
- van den Bergh (1986) van den Bergh, S. 1986, AJ, 91, 271
- van Dokkum et al. (2015) van Dokkum, P. G., Abraham, R., Merritt, A., et al. 2015, ApJ, 798, L45
- Voggel et al. (2016) Voggel, K., Hilker, M., & Richtler, T. 2016, A&A, 586, A102
- Wang et al. (2023) Wang, K., Peng, E. W., Liu, C., et al. 2023, Nature, 623, 296
- Wittmann et al. (2016) Wittmann, C., Lisker, T., Pasquali, A., Hilker, M., & Grebel, E. K. 2016, MNRAS, 459, 4450
- Zanatta et al. (2021) Zanatta, E., Sánchez-Janssen, R., Chies-Santos, A. L., de Souza, R. S., & Blakeslee, J. P. 2021, MNRAS, 508, 986
- Zinnecker et al. (1988) Zinnecker, H., Keable, C. J., Dunlop, J. S., Cannon, R. D., & Griffiths, W. K. 1988, in IAU Symposium, Vol. 126, The Harlow-Shapley Symposium on Globular Cluster Systems in Galaxies, ed. J. E. Grindlay & A. G. D. Philip, 603
Appendix A NSCs properties
Table 1 presents the structural and photometric properties of the 43 NSCs identified in the 41 nucleated dwarf galaxies observed with HST. For each NSC, the table includes the host galaxy ID, distance and stellar mass (M), together with the NSC RA/Dec coordinates, extinction corrected absolute magnitude in the -band (MV,NSC), color, effective radius (Re,NSC), Sérsic index , ellipticity, and stellar mass (M). Two values of M are shown: the one derived for this work based on Roediger & Courteau (2015) stellar mass-to-light ratios, and in parentheses, the one used in either Poulain et al. (2021) or Poulain et al. (2022) based on Bell et al. (2003) method. The last column indicates the band from which the parameters were derived. If the parameters are based on one band, 2Re,NSC aperture photometry was applied in both bands to derive the color. When both bands were used, Re,NSC, , and the ellipticity correspond to the average value of both bands. An absence of band means that there is no reliable model for the NSC, and that the derived photometry was picked from either Poulain et al. (2025) or Marleau et al. (2024), when available. The NSCs are sorted by increasing RA. MATLAS-138 and MATLAS-987 have a double nucleus, and thus have two lines.
@@ Appendix A NSCs properties
Appendix B NSCs color profiles

### v2

@@ 9 Summary and conclusions
In this work we investigated the formation and evolution of NSCs by studying their structural and photometric properties in a dwarf galaxy sample biased toward low surface brightness and diffuse (UDG-like) galaxies located beyond the LV and outside the environment of galaxy clusters. We made use of HST follow-up observations of 79 MATLAS galaxies with the ACS using the F606W and F814W filters. This sample is composed of 41 nucleated dwarfs, 13 of which are newly identified thanks to HST high spatial resolution images, and two have a double nucleus.
@@
We modeled the NSCs with a Sérsic profile and obtained a reliable result for 31 of them; the derived parameters are provided in Table 1. We compared their structural properties and photometry to NSCs in nearby field LTGs, and in dwarf galaxies from the LV, and from the Virgo and Fornax galaxy clusters. The Re,NSC of the MATLAS NSCs does not tend to be larger as they get brighter and more massive. Moreover, while we do not see an influence of the absolute magnitude on the ellipticity, the Sérsic index of the MATLAS NSCs gets larger with M, as observed in the Virgo cluster.
@@
The relation between the NSCs and their host galaxy shows that a non-negligible fraction of the MATLAS NSCs tend to be brighter and more massive than NSCs in typical dwarf galaxies. This could be due to the galaxies undergoing tidal disruption or to an increased growth of NSCs in UDGs, thus producing more massive NSCs.
@@
We derived the color profiles within 3Re,NSC of the modeled NSCs and remark that, while the low-mass NSCs tend to have an almost constant profile consistent with the migration scenario, the most massive NSCs have a steep slope, from an outer redder color to an inner bluer color, suggesting a change in stellar population, which might be caused by a contribution from in situ star formation.
@@
Following the work of Marleau2024 on the GC systems of the HST MATLAS UDGs, we studied the distribution and properties of the GCs in the context of the migration scenario. We modeled the GCs in the MATLAS dwarfs up to 25 Mpc and find that, although with low statistics, the most massive GCs with M¿ 10M⊙ are distributed oppositely in the nucleated and non-nucleated dwarfs: the former seem to lack massive GCs in their central 0.5Re,gal region, while they are located within 0.5Re,gal for the latter. Comparing the Re – luminosity scaling relation of the GCs to the NSCs, we find a different relation for the two types of star clusters, in agreement with Bekki2004.
@@
The most massive UCDs are believed to be the NSC remnants of tidally stripped nucleated dwarfs. In that context, we searched for UCD progenitors among the 41 NSCs with available photometry. We found that the NSCs appear to link the GC and UCD populations in the Re – luminosity scaling relation, suggesting that some of them could become UCDs if their dwarf host undergo tidal disruption. Looking for strongly nucleated dwarfs, we identified two UCD progenitor candidates, MATLAS-1539 and MATLAS-1577, that have a nuclear-to-total-luminosity fraction of 6.7% and 8.7%, respectively, and show signs of a possible tidal interaction with their assumed host ETG. For the galaxies with a low nuclear-to-total-luminosity fraction, we found three tidally elongated UDGs, likely in the early stage of tidal disruption, which we added to our list of UCD progenitor candidates.
@@
Overall, the observed trends need to be confirmed by studying even larger samples of nucleated dwarfs using a deep and high-resolution imaging facility (e.g., the Euclid Wide Survey; Scaramella2022), combined with spectroscopic follow-up observations to trace the NSC star formation history.
@@ Acknowledgements.
The authors would like to thank the referee for the insightful comments, which helped to improve the manuscript. This research is based on observations from the NASA/ESA Hubble Space Telescope obtained at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. Support for Program number GO-16257 and GO-16711 was provided through a grant from the STScI under NASA contract NAS5-26555. This research was supported by the International Space Science Institute (ISSI) in Bern, through ISSI International Team project No. 534. This research made use of AutoProf, a package for galaxy image photometry (Stone2021b). M.P. is supported by the Academy of Finland grant No. 347089. O.M. and N.H. are grateful to the Swiss National Science Foundation for financial support under the grant number PZ00P2_202104. P.R.D. gratefully acknowledges support from grant HST-GO-16257.002-A. S.P. acknowledges support from the Mid-career Researcher Program (No. RS-2023-00208957). S.L. acknowledges the support from the Sejong Science Fellowship Program by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. NRF-2021R1C1C2006790). RH acknowledges funding from the Italian INAF Large Grant 12 - 2022.Appendix A NSC properties
Table 1 presents the structural and photometric properties of the 43 NSCs identified in the 41 nucleated dwarf galaxies observed with HST. For each NSC, the table includes the host galaxy ID, distance and stellar mass (M), together with the NSC RA/Dec coordinates, extinction corrected absolute magnitude in the -band (MV,NSC), color, effective radius (Re,NSC), Sérsic index , ellipticity, and stellar mass (M). Two values of M are shown: the one derived for this work based on Roediger2015 stellar mass-to-light ratios, and in parentheses, the one used in either Poulain2021 or Poulain2022 based on Bell2003 method. The last column indicates the band from which the parameters were derived. If the parameters are based on one band, 2Re,NSC aperture photometry was applied in both bands to derive the color. When both bands were used, Re,NSC, , and the ellipticity correspond to the average value of both bands. An absence of band means that there is no reliable model for the NSC, and that the derived photometry was picked from either Poulain2025 or Marleau2024, when available. The NSCs are sorted by increasing RA. MATLAS-138 and MATLAS-987 have a double nucleus, and thus have two lines.
@@
∗ NSC Re smaller than one pixel.
Distance measurement from aAnn2015, bheesters2023, and cSDSS DR13 database (Albareti2017; Blanton2017).
Appendix B NSC color profiles